1) If the total utilities for the fifth and sixth units of a good consumed are 98 and 115 respectively. The marginal utility for the sixth unit is:


1
Expert's answer
2021-04-30T10:47:16-0400

Marginal utility is the change in the utility from the increase in consumption of goods. It is the added satisfaction of the consumer for consuming one more unit of good or service. The marginal utility can be either positive, negative or zero.



Marginal utility,(MU)"=\frac{\Delta TU}{\Delta Q}"

where,

∆TU = change in total utility∆Q =change in number of units consumed
MU for the 6th unit "=\frac{115-98}{6-5}=17"

Therefore marginal utility for the sixth unit is 17